# SpanTrain: Highly Efficient Cross-domain Model Distributed Training System under Heterogeneous GPUs and Networks in CEE Environment

### v1

@@ Abstract.
Most existing training systems focus on a single region. In contrast, we envision that cross-region training offers more flexible GPU resource allocation and yields significant potential. However, the hierarchical cluster topology and unstable networks in the cloud-edge-end (CEE) environment, a typical cross-region scenario, pose substantial challenges to building an efficient and autonomous model training system. We propose SpanTrain, a geo-distributed model training system tailored for heterogeneous GPUs and networks in CEE environments. SpanTrain adopts a communication-centric design philosophy to tackle challenges arising from slow and unstable inter-region networks. It begins with a heterogeneous device profiler that identifies and groups devices based on both network and compute characteristics. Leveraging device groups, SpanTrain implements compact, zero-bubble pipeline parallelism, automatically deriving optimal parallel strategies. To further adapt to runtime variability, SpanTrain integrates a dynamic environment adapter that reacts to network fluctuations. Extensive evaluations demonstrate that SpanTrain achieves 1.3-2.8$\times$ higher training throughput compared to widely used and SOTA training systems.
@@ 1. Introduction
We present SpanTrain (§3), a geo-distributed model training system that can automatically generate parallel strategies and adapt to network fluctuations in the CEE environment. The core idea of SpanTrain is network-centric automatic parallel strategy generation and adjustment. While most existing work focuses on the impact of compute heterogeneity on parallel strategies, in CEE environments, network heterogeneity becomes the primary factor affecting training efficiency. This makes it difficult for existing approaches to fully utilize GPU compute resources across regions. SpanTrain introduces a network-centric approach, prioritizing the mitigation of pipeline bubbles caused by cross-region networks in device grouping, automatic parallel strategy generation, and the adjustment of pipeline stages during execution. Specifically, SpanTrain includes the following three key techniques.
First, SpanTrain introduces the Heterogeneous Devices Profiler, which plans a two-level device grouping before training begins. The first-level groups GPUs based on network performance, eliminating the impact of cross-region network hierarchies on parallel training strategies. The second level further groups GPUs within each network-isolated group based on compute capability, clustering GPUs with similar computation and memory access performance.
Second, SpanTrain proposes the Parallel Strategy Planner, which adopts different parallel strategies for each level of the two-level device group. At the first level, considering the low bandwidth and variability of cross-region networks, SpanTrain introduces compact zero-bubble pipeline parallelism, where weight computations during back propagation are decoupled and inserted into potential pipeline bubbles in a compact fashion. At the second level, SpanTrain automatically generates the optimal combination of pipeline, data and tensor parallelism by jointly considering differences in compute and memory access capabilities across GPUs.
@@
Third, SpanTrain introduces the Dynamic Environment Adapter to cope with network fluctuations. Unlike existing methods that rely on checkpoints and pipeline restarts, SpanTrain monitors network variability and adapts the system through micro-batch-level adjustments, enabling continuous and efficient training without interruptions.
We implement SpanTrain based on PyTorch with around 15K lines of code. We evaluate SpanTrain against a wide range of state-of-the-art training systems including PyTorch DDP (Li et al., 2020), PyTorch FSDP (Zhao et al., 2023), DeepSpeed ZeRO2 (Rajbhandari et al., 2020), DeepSpeed Pipeline (Narayanan et al., 2019), Gpipe (Huang et al., 2019), Alpa (Zheng et al., 2022), Metis (Um et al., 2024), Hetpipe (Um et al., 2024), and Asteroid (Ye et al., 2024). Extensive results (§4) in a simulated CEE environment show that the SpanTrain outperforms existing distributed training systems by 1.3-2.8$\times$ in terms of training throughput. Furthermore, under dynamic network changes, SpanTrain improves training throughput by 1.5-1.7$\times$ compared to existing systems.
@@
• We propose SpanTrain, a geo-distributed training system designed specifically for CEE computing environments, with a network-centric design philosophy at its core.
• We implement SpanTrain based on PyTorch and conduct extensive evaluations to verify its effectiveness.
@@ 2.1. Challenges 1: Hierarchical Cluster Topology
Figure 4. The main working components and workflow of SpanTrain. SpanTrain includes the pre-run performance evaluation component Heterogeneous Devices Profiler, the pre-run parallel planning component Parallel Strategy Planner, and the runtime environment adaptation component Dynamic Environment Adapter.
@@
The main working components and workflow of SpanTrain. SpanTrain includes the pre-run performance evaluation component Heterogeneous Devices Profiler, the pre-run parallel plan planning component Parallel Strategy Planner, and the runtime environment adaptation component Dynamic Environment Adapter.
@@ 2.2. Challenges 2: Network Fluctuations
3. SpanTrain: Design and Implementation
We present SpanTrain, a cross-region model training system designed to address the unique challenges of the CEE environment, as discussed in §2. This section first provides an overview (§ 3.1) and highlights the core idea behind SpanTrain, then delves into the key techniques, namely Heterogeneous Devices Profiler (§ 3.2), Parallel Strategy Planner (§ 3.3), and Dynamic Environment Adapter (§ 3.4).
@@ 3.1. Key Idea and Overview
Departing from the compute-centric approach of existing model training systems, SpanTrain generates and adapts parallelization strategies driven by network performance. In the CEE environment, where the cluster topology exhibits a hierarchical characteristic in terms of computing device and network, in addition, network performance also varies significantly across different tiers and exhibits noticeable fluctuations (as shown in §2). Existing systems often mistakenly group cross-region devices into a single device group, resulting in suboptimal parallelization strategies and increased pipeline bubbles during network fluctuations. By treating network performance evaluation as a first-class citizen, SpanTrain effectively addresses the network challenges inherent in cross-region training, with employing computing device performance evaluation as a second-class citizen.
Figure 4 presents the overview and workflow of SpanTrain. SpanTrain first employs the Profiler (§3.2) to find optimal device groups. It then uses the Planner (§3.3) to automatically generate an asymmetric hybrid parallel strategy. Finally, during periods of network fluctuation, it introduces Adapter (§3.4) to mitigate pipeline bubbles, thereby ensuring stable training performance. Specifically, the core components of SpanTrain are as follows:
@@
• Parallel Strategy Planner (§ 3.3). It introduces an asymmetric hybrid parallel mechanism based on the two-level device group. Among various training parallelism strategies, pipeline parallelism offers the lowest network overhead, making it the preferred choice for communication across the first-level device groups in SpanTrain. Given the slow cross-region networks, conventional 1F1B (one-forward-one-backward) pipeline execution introduces significant pipeline bubbles, resulting in suboptimal compute utilization. To alleviate this inefficiency, SpanTrain adopts zero pipeline (Qi et al., 2023), which separates weight computations during backpropagation and fills idle execution slots, thereby improving compute resource utilization under slow network conditions. In the second-level device group, SpanTrain autonomously derives a 3D parallelism strategy guided by the principle of load balancing across heterogeneous compute resources. This involves sequentially searching for the optimal combination of pipeline, data, and tensor parallelism.
@@ 3.3. Parallel Strategy Planner
Conventional parallelism strategies are ineffective for CEE environments (§2.1). To address this, SpanTrain introduces the Planner, which implements a compact zero-bubble pipeline architecture that decouples weight updates from gradient computation and proactively inserts the weight updates to pipeline bubbles incurred by the slow and heterogeneous networks whenever feasible. Building upon this foundation, SpanTrain further integrates a region-aware cost model alongside an automated strategy search algorithm to collaboratively generate optimized parallel execution plans.
@@
The pipeline architecture and cost modeling stage of SpanTrain. (G:first-level device group)
@@ 3.3.1. Compact Zero-bubble Pipeline Architecture
As that pipeline parallelism (PP) incurs the lowest communication overhead, SpanTrain adopts PP across first-level groups. In particular, the Planner implements compact zero-bubble PP, building upon the original zero-bubble pipeline design (Qi et al., 2023), which decouples gradient computation from weight updates. This decoupling substantially enlarges the pipeline scheduling space. However, the original zero-bubble PP assumes that inter-stage communication is faster than both forward and backward propagation computation. This leads to a rigid execution pattern where forward passes, backward passes, and weight updates follow fixed, non-overlapping phases. Original zero-bubble pipeline generally follows the 1F1B schedule, but it adjusts the starting points of weight update depending on the number of warm-up micro-sbatches. However, CEE’s hierarchical cluster topology and bandwidth-constrained networks violate its fundamental assumption. As Figure 6a demonstrates, applying the original zero-bubble pipeline to CEE environments exposes critical limitations. Its fixed scheduling pattern generates excessive GPU bubbles that leave most of computing resources idle during network bandwidth-constrained periods.
@@ 3.3.2. Cost Model
To find optimal parallel strategies for CEE environments, SpanTrain first formalizes their costs by Equation 1. It then tries to find a solution that minimizes the following cost ($T^{*}$) based on the compact zero-bubble PP.
@@ 4. Evaluation
We will answer the following questions experimentally: What and why does SpanTrain improve performance compared to the state-of-the-art (SOTA) works in the CEE when training diverse models (§4.2) ? What is the effectiveness and adaptation of SpanTrain’s parallel strategy compared to the SOTA works (§4.3 and §4.4)? What is the scalability of SpanTrain compared to the SOTA works (§4.5) ? What is the SpanTrain’s system overhead (§4.6) ?
@@ 4.1. Evaluation Environment
Models and Datasets. We evaluate SpanTrain with 5 typical AI models that are widely used in the CEE, including CV models (i.e., ResNet50, EfficientNet) and NLP models (i.e., Bert-small, GPT-2). We use Cifar-10 dataset with input $3\times 32\times 32$ for ResNet101, EfficientNet. For Bert-small, GPT-2, we use the oscar-en-10k dataset with input shape $32\times 512$.
@@
Compared Methods. We compare SpanTrain with both widely-used frameworks and the SOTA methods:
@@ 4.2. End-to-End Performance Evaluation
This section evaluates the training throughput and time overhead of SpanTrain against widely-used training frameworks in achieving the specified model training accuracy.
@@
Training Throughput. As shown in Figure 9b, SpanTrain improves training performance by 2.3-2.38$\times$ and 1.74-1.85$\times$ compared to PyTorch DDP in settings 1 and 2, respectively. It achieves 2.05–2.11$\times$ and 1.61–1.76$\times$ performance gains compared to PyTorch FSDP in settings 1 and 2, respectively. Similarly, SpanTrain delivers 1.04–1.57$\times$ and 1.13–1.71$\times$ performance improvements compared to DeepSpeed Pipeline and GPipe, respectively. These performance gains are primarily due to SpanTrain ’s ability to effectively address device heterogeneity and dynamic network diversity in CEE environments, which are overlooked by both PyTorch DDP, PyTorch FDSP, DeepSpeed Pipeline and Gpipe, resulting in worker imbalance and inefficient synchronization. We further observe an intriguing phenomenon in the CEE: larger batch sizes consistently deliver higher throughput. Specifically, our evaluations demonstrate a 6.8-16.2% throughput improvement when increasing batch size from 128 to 256 (Figure 9b). This stems from two key factors: (1) longer computing intervals create more opportunities for communication-computation overlap, (2) increased micro-batch counts compress pipeline warm-up phases.
@@
Time Overhead. As shown in Figure 10d, the time overhead is evaluated with using different models when achieving the same accuracy in Setting 1. Specifically, training terminates when ResNet-101 and EfficientNet achieve the target accuracy of 85%, and when Bert-small and GPT-2 achieve the target BLEU of 0.65. Compared with existing approaches, SpanTrain takes 55-73.5% less time. This is mainly thanks to SpanTrain can adapt to the CEE by a reasonable hybrid parallel training strategy.
@@
However, although SpanTrain can get better training performance than existing methods, it does not fully translate the training throughput improvement into training acceleration, resulting in the training time increased by 6.4-8.2% in contrast to theoretical improvement. This occurs because our compact zero-bubble pipeline disrupts the conventional weight update sequence, forward and backward no longer strictly operate on the recent model parameters. While such aggressive out-of-order execution improves average throughput by 68.1% (Figure 9b), it introduces parameter instability that requires 6.4-8.2% more training iterations to reach target accuracy (Figure 10d). Nevertheless, SpanTrain still achieves faster end-to-end training compared to other approaches, making this trade-off acceptable for most deployments.
4.3. Effectiveness Evaluation of SpanTrain
@@
This section compares SpanTrain with SOTA frameworks, in terms of training throughput and search cost of parallel strategies. We use Setting 1 for evaluation.
@@
4.3.1  Training Throughput. As shown in Figure 11a, when evaluating the training throughput of ResNet, EfficientNet, Bert-small, and GPT-2 using different auto-parallel training frameworks, SpanTrain can outperform existing frameworks in most cases by 1.49-2.85$\times$.
@@
Comparison against Alpa. SpanTrain offers 2.46–2.85$\times$ higher training throughput compared to Alpa. While Alpa pioneered rapid parallel strategy search through unified inter-op/intra-op parallelism, combining tensor, data, and pipeline parallelism via integer linear programming. However, its optimization targets homogeneous clusters (e.g., uniform 8-GPU nodes) and assumes static hardware conditions. This leads to rigid, evenly partitioned model sharding that ignores device heterogeneity and network variance. In contrast, SpanTrain incorporates Profiler that obtains device compute capabilities and network bandwidth, enabling dynamic load balancing adapted to real-time hardware performance.
@@
Comparison against Metis. Our experiments also compare SpanTrain with Metis, a variant of Alpa. While Metis demonstrates measurable throughput improvements over Alpa in our test environment by accounting for device heterogeneity, SpanTrain maintains a 1.49-1.52$\times$ throughput advantage. This performance gap stems from Metis’ simplified heterogeneity model, which only addresses basic inter-cluster differences between two heterogeneous computing groups while neglecting network diversity and hierarchy. Metis’ sole optimization objective of minimizing inter-cluster communication consequently yields limited performance gains.
@@
Comparison against HetPipe. SpanTrain demonstrates superior training throughput, achieving 2.03-2.18$\times$ improvements over HetPipe. HetPipe coordinates heterogeneous devices and diverse networks through communication scheduling. But, its parameter server (PS) architecture necessitates epoch-level weight synchronization that halts training and limits throughput. However, SpanTrain’s fully asynchronous pipeline parallelism eliminates this bottleneck.
@@
Comparsion against Asteroid. SpanTrain achieves 1.04-1.05$\times$ higher training throughput than Asteroid. Asteroid uses dynamic programming based multi-dimensional search approach to achieve similar performance. However, SpanTrain’s strategy more comprehensively accounts for the CEE hierarchical structures, avoiding the suboptimal parallel strategy generated by Asteroid’s overcomplexed assumptions. Furthermore, SpanTrain’s significantly faster strategy search enables quicker training initialization, and its dynamic microbatch scheduling combined with asynchronous execution provides additional performance gains beyond Asteroid’s capabilities.
@@
4.3.2  Cost Time. As illustrated in Figure 11b, Alpa, Metis, and Asteroid exhibit substantially longer parallelism strategy search time compared to SpanTrain. Alpa requires 286.3 seconds for ResNet101 and 130 seconds for GPT-2, primarily due to its exhaustive search across all possible device combinations, layer mappings, and hyperparameters (e.g., batch/micro-batch sizes), resulting in an $O(n!\cdot n!\cdot n^{2})$ search space. Metis reduces this overhead to 230 seconds (ResNet101) and 43 seconds (GPT-2) by pre-grouping devices into performance, balanced device groups and applying search-space pruning, though it still maintains $O(n!\cdot n^{3})$ time complexity. Asteroid further shortens the search time to 197 seconds (ResNet101) and 59 seconds (GPT-2) but remains slower than SpanTrain because it (1) operates on the full device set without pre-grouping and (2) incurs additional latency from runtime memory analysis and parallel strategy evaluation.
In contrast, SpanTrain achieves significantly faster strategy discovery through three key optimizations: (1) topology-aware device grouping that reduces the search space to near-constant complexity, (2) rapid candidate generation via predefined partitioning points, and (3) lightweight refinement using beam search.
4.3.3  Profit of Second-Level Computing Device Group. In Figure 11a, we evaluate the training performance when turning off SpanTrain’s model split among second-level device groups (No-Second). The experimental results show that the training performance of No-Second declines by 10.93% and 12.04% compared to SpanTrain when using ResNet101 and GPT-2, respectively. This is due to the inability to fully leverage the resources of the heterogeneous computing devices when model splitting is disabled among second-level device groups. However, its training performance is still higher than existing parallel training frameworks such as Alpa, Metis, and Hetpipe. As shown in Figure 11b, turning off SpanTrain’s second-level device group search can reduce its search time of the parallel strategy by roughly 20%.
4.4. Evaluating SpanTrain’s Adaptation

### v2

# DeepCEE: Efficient Cross-Region Model Distributed Training System under Heterogeneous GPUs and Networks

@@ Abstract.
Most existing training systems focus on a single region. In contrast, we envision that cross-region training offers more flexible GPU resource allocation and yields significant potential. However, the hierarchical cluster topology and unstable networks in the cloud-edge-end (CEE) environment, a typical cross-region scenario, pose substantial challenges to building an efficient and autonomous model training system. We propose DeepCEE, a geo-distributed model training system tailored for heterogeneous GPUs and networks in CEE environments. DeepCEE adopts a communication-centric design philosophy to tackle challenges arising from slow and unstable inter-region networks. It begins with a heterogeneous device profiler that identifies and groups devices based on both network and compute characteristics. Leveraging device groups, DeepCEE implements compact, zero-bubble pipeline parallelism, automatically deriving optimal parallel strategies. To further adapt to runtime variability, DeepCEE integrates a dynamic environment adapter that reacts to network fluctuations. Extensive evaluations demonstrate that DeepCEE achieves 1.3-2.8$\times$ higher training throughput compared to widely used and SOTA training systems.
@@ 1. Introduction
We present DeepCEE (§3), a geo-distributed model training system that can automatically generate parallel strategies and adapt to network fluctuations in the CEE environment. The core idea of DeepCEE is network-centric automatic parallel strategy generation and adjustment. While most existing work focuses on the impact of compute heterogeneity on parallel strategies, in CEE environments, network heterogeneity becomes the primary factor affecting training efficiency. This makes it difficult for existing approaches to fully utilize GPU compute resources across regions. DeepCEE introduces a network-centric approach, prioritizing the mitigation of pipeline bubbles caused by cross-region networks in device grouping, automatic parallel strategy generation, and the adjustment of pipeline stages during execution. Specifically, DeepCEE includes the following three key techniques.
First, DeepCEE introduces the Heterogeneous Devices Profiler, which plans a two-level device grouping before training begins. The first-level groups GPUs based on network performance, eliminating the impact of cross-region network hierarchies on parallel training strategies. The second level further groups GPUs within each network-isolated group based on compute capability, clustering GPUs with similar computation and memory access performance.
Second, DeepCEE proposes the Parallel Strategy Planner, which adopts different parallel strategies for each level of the two-level device group. At the first level, considering the low bandwidth and variability of cross-region networks, DeepCEE introduces compact zero-bubble pipeline parallelism, where weight computations during back propagation are decoupled and inserted into potential pipeline bubbles in a compact fashion. At the second level, DeepCEE automatically generates the optimal combination of pipeline, data and tensor parallelism by jointly considering differences in compute and memory access capabilities across GPUs.
@@
Third, DeepCEE introduces the Dynamic Environment Adapter to cope with network fluctuations. Unlike existing methods that rely on checkpoints and pipeline restarts, DeepCEE monitors network variability and adapts the system through micro-batch-level adjustments, enabling continuous and efficient training without interruptions.
We implement DeepCEE based on PyTorch with around 15K lines of code. We evaluate DeepCEE against a wide range of state-of-the-art training systems including PyTorch DDP (Li et al., 2020), PyTorch FSDP (Zhao et al., 2023), DeepSpeed ZeRO2 (Rajbhandari et al., 2020), DeepSpeed Pipeline (Narayanan et al., 2019), Gpipe (Huang et al., 2019), Alpa (Zheng et al., 2022), Metis (Um et al., 2024), Hetpipe (Um et al., 2024), and Asteroid (Ye et al., 2024). Extensive results (§4) in a simulated CEE environment show that the DeepCEE outperforms existing distributed training systems by 1.3-2.8$\times$ in terms of training throughput. Furthermore, under dynamic network changes, DeepCEE improves training throughput by 1.5-1.7$\times$ compared to existing systems.
@@
• We propose DeepCEE, a geo-distributed training system designed specifically for CEE computing environments, with a network-centric design philosophy at its core.
• We implement DeepCEE based on PyTorch and conduct extensive evaluations to verify its effectiveness.
@@ 2.1. Challenges 1: Hierarchical Cluster Topology
Figure 4. The main working components and workflow of DeepCEE. DeepCEE includes the pre-run performance evaluation component Heterogeneous Devices Profiler, the pre-run parallel planning component Parallel Strategy Planner, and the runtime environment adaptation component Dynamic Environment Adapter.
@@
The main working components and workflow of DeepCEE. DeepCEE includes the pre-run performance evaluation component Heterogeneous Devices Profiler, the pre-run parallel plan planning component Parallel Strategy Planner, and the runtime environment adaptation component Dynamic Environment Adapter.
@@ 2.2. Challenges 2: Network Fluctuations
3. DeepCEE: Design and Implementation
We present DeepCEE, a cross-region model training system designed to address the unique challenges of the CEE environment, as discussed in §2. This section first provides an overview (§ 3.1) and highlights the core idea behind DeepCEE, then delves into the key techniques, namely Heterogeneous Devices Profiler (§ 3.2), Parallel Strategy Planner (§ 3.3), and Dynamic Environment Adapter (§ 3.4).
@@ 3.1. Key Idea and Overview
Departing from the compute-centric approach of existing model training systems, DeepCEE generates and adapts parallelization strategies driven by network performance. In the CEE environment, where the cluster topology exhibits a hierarchical characteristic in terms of computing device and network, in addition, network performance also varies significantly across different tiers and exhibits noticeable fluctuations (as shown in §2). Existing systems often mistakenly group cross-region devices into a single device group, resulting in suboptimal parallelization strategies and increased pipeline bubbles during network fluctuations. By treating network performance evaluation as a first-class citizen, DeepCEE effectively addresses the network challenges inherent in cross-region training, with employing computing device performance evaluation as a second-class citizen.
Figure 4 presents the overview and workflow of DeepCEE. DeepCEE first employs the Profiler (§3.2) to find optimal device groups. It then uses the Planner (§3.3) to automatically generate an asymmetric hybrid parallel strategy. Finally, during periods of network fluctuation, it introduces Adapter (§3.4) to mitigate pipeline bubbles, thereby ensuring stable training performance. Specifically, the core components of DeepCEE are as follows:
@@
• Parallel Strategy Planner (§ 3.3). It introduces an asymmetric hybrid parallel mechanism based on the two-level device group. Among various training parallelism strategies, pipeline parallelism offers the lowest network overhead, making it the preferred choice for communication across the first-level device groups in DeepCEE. Given the slow cross-region networks, conventional 1F1B (one-forward-one-backward) pipeline execution introduces significant pipeline bubbles, resulting in suboptimal compute utilization. To alleviate this inefficiency, DeepCEE adopts zero pipeline (Qi et al., 2023), which separates weight computations during backpropagation and fills idle execution slots, thereby improving compute resource utilization under slow network conditions. In the second-level device group, DeepCEE autonomously derives a 3D parallelism strategy guided by the principle of load balancing across heterogeneous compute resources. This involves sequentially searching for the optimal combination of pipeline, data, and tensor parallelism.
@@ 3.3. Parallel Strategy Planner
Conventional parallelism strategies are ineffective for CEE environments (§2.1). To address this, DeepCEE introduces the Planner, which implements a compact zero-bubble pipeline architecture that decouples weight updates from gradient computation and proactively inserts the weight updates to pipeline bubbles incurred by the slow and heterogeneous networks whenever feasible. Building upon this foundation, DeepCEE further integrates a region-aware cost model alongside an automated strategy search algorithm to collaboratively generate optimized parallel execution plans.
@@
The pipeline architecture and cost modeling stage of DeepCEE. (G:first-level device group)
@@ 3.3.1. Compact Zero-bubble Pipeline Architecture
As that pipeline parallelism (PP) incurs the lowest communication overhead, DeepCEE adopts PP across first-level groups. In particular, the Planner implements compact zero-bubble PP, building upon the original zero-bubble pipeline design (Qi et al., 2023), which decouples gradient computation from weight updates. This decoupling substantially enlarges the pipeline scheduling space. However, the original zero-bubble PP assumes that inter-stage communication is faster than both forward and backward propagation computation. This leads to a rigid execution pattern where forward passes, backward passes, and weight updates follow fixed, non-overlapping phases. Original zero-bubble pipeline generally follows the 1F1B schedule, but it adjusts the starting points of weight update depending on the number of warm-up micro-sbatches. However, CEE’s hierarchical cluster topology and bandwidth-constrained networks violate its fundamental assumption. As Figure 6a demonstrates, applying the original zero-bubble pipeline to CEE environments exposes critical limitations. Its fixed scheduling pattern generates excessive GPU bubbles that leave most of computing resources idle during network bandwidth-constrained periods.
@@ 3.3.2. Cost Model
To find optimal parallel strategies for CEE environments, DeepCEE first formalizes their costs by Equation 1. It then tries to find a solution that minimizes the following cost ($T^{*}$) based on the compact zero-bubble PP.
@@ 4. Evaluation
We will answer the following questions experimentally: What and why does DeepCEE improve performance compared to the state-of-the-art (SOTA) works in the CEE when training diverse models (§4.2) ? What is the effectiveness and adaptation of DeepCEE’s parallel strategy compared to the SOTA works (§4.3 and §4.4)? What is the scalability of DeepCEE compared to the SOTA works (§4.5) ? What is the DeepCEE’s system overhead (§4.6) ?
@@ 4.1. Evaluation Environment
Models and Datasets. We evaluate DeepCEE with 5 typical AI models that are widely used in the CEE, including CV models (i.e., ResNet50, EfficientNet) and NLP models (i.e., Bert-small, GPT-2). We use Cifar-10 dataset with input $3\times 32\times 32$ for ResNet101, EfficientNet. For Bert-small, GPT-2, we use the oscar-en-10k dataset with input shape $32\times 512$.
@@
Compared Methods. We compare DeepCEE with both widely-used frameworks and the SOTA methods:
@@ 4.2. End-to-End Performance Evaluation
This section evaluates the training throughput and time overhead of DeepCEE against widely-used training frameworks in achieving the specified model training accuracy.
@@
Training Throughput. As shown in Figure 9b, DeepCEE improves training performance by 2.3-2.38$\times$ and 1.74-1.85$\times$ compared to PyTorch DDP in settings 1 and 2, respectively. It achieves 2.05–2.11$\times$ and 1.61–1.76$\times$ performance gains compared to PyTorch FSDP in settings 1 and 2, respectively. Similarly, DeepCEE delivers 1.04–1.57$\times$ and 1.13–1.71$\times$ performance improvements compared to DeepSpeed Pipeline and GPipe, respectively. These performance gains are primarily due to DeepCEE ’s ability to effectively address device heterogeneity and dynamic network diversity in CEE environments, which are overlooked by both PyTorch DDP, PyTorch FDSP, DeepSpeed Pipeline and Gpipe, resulting in worker imbalance and inefficient synchronization. We further observe an intriguing phenomenon in the CEE: larger batch sizes consistently deliver higher throughput. Specifically, our evaluations demonstrate a 6.8-16.2% throughput improvement when increasing batch size from 128 to 256 (Figure 9b). This stems from two key factors: (1) longer computing intervals create more opportunities for communication-computation overlap, (2) increased micro-batch counts compress pipeline warm-up phases.
@@
Time Overhead. As shown in Figure 10d, the time overhead is evaluated with using different models when achieving the same accuracy in Setting 1. Specifically, training terminates when ResNet-101 and EfficientNet achieve the target accuracy of 85%, and when Bert-small and GPT-2 achieve the target BLEU of 0.65. Compared with existing approaches, DeepCEE takes 55-73.5% less time. This is mainly thanks to DeepCEE can adapt to the CEE by a reasonable hybrid parallel training strategy.
@@
However, although DeepCEE can get better training performance than existing methods, it does not fully translate the training throughput improvement into training acceleration, resulting in the training time increased by 6.4-8.2% in contrast to theoretical improvement. This occurs because our compact zero-bubble pipeline disrupts the conventional weight update sequence, forward and backward no longer strictly operate on the recent model parameters. While such aggressive out-of-order execution improves average throughput by 68.1% (Figure 9b), it introduces parameter instability that requires 6.4-8.2% more training iterations to reach target accuracy (Figure 10d). Nevertheless, DeepCEE still achieves faster end-to-end training compared to other approaches, making this trade-off acceptable for most deployments.
4.3. Effectiveness Evaluation of DeepCEE
@@
This section compares DeepCEE with SOTA frameworks, in terms of training throughput and search cost of parallel strategies. We use Setting 1 for evaluation.
@@
4.3.1  Training Throughput. As shown in Figure 11a, when evaluating the training throughput of ResNet, EfficientNet, Bert-small, and GPT-2 using different auto-parallel training frameworks, DeepCEE can outperform existing frameworks in most cases by 1.49-2.85$\times$.
@@
Comparison against Alpa. DeepCEE offers 2.46–2.85$\times$ higher training throughput compared to Alpa. While Alpa pioneered rapid parallel strategy search through unified inter-op/intra-op parallelism, combining tensor, data, and pipeline parallelism via integer linear programming. However, its optimization targets homogeneous clusters (e.g., uniform 8-GPU nodes) and assumes static hardware conditions. This leads to rigid, evenly partitioned model sharding that ignores device heterogeneity and network variance. In contrast, DeepCEE incorporates Profiler that obtains device compute capabilities and network bandwidth, enabling dynamic load balancing adapted to real-time hardware performance.
@@
Comparison against Metis. Our experiments also compare DeepCEE with Metis, a variant of Alpa. While Metis demonstrates measurable throughput improvements over Alpa in our test environment by accounting for device heterogeneity, DeepCEE maintains a 1.49-1.52$\times$ throughput advantage. This performance gap stems from Metis’ simplified heterogeneity model, which only addresses basic inter-cluster differences between two heterogeneous computing groups while neglecting network diversity and hierarchy. Metis’ sole optimization objective of minimizing inter-cluster communication consequently yields limited performance gains.
@@
Comparison against HetPipe. DeepCEE demonstrates superior training throughput, achieving 2.03-2.18$\times$ improvements over HetPipe. HetPipe coordinates heterogeneous devices and diverse networks through communication scheduling. But, its parameter server (PS) architecture necessitates epoch-level weight synchronization that halts training and limits throughput. However, DeepCEE’s fully asynchronous pipeline parallelism eliminates this bottleneck.
@@
Comparsion against Asteroid. DeepCEE achieves 1.04-1.05$\times$ higher training throughput than Asteroid. Asteroid uses dynamic programming based multi-dimensional search approach to achieve similar performance. However, DeepCEE’s strategy more comprehensively accounts for the CEE hierarchical structures, avoiding the suboptimal parallel strategy generated by Asteroid’s overcomplexed assumptions. Furthermore, DeepCEE’s significantly faster strategy search enables quicker training initialization, and its dynamic microbatch scheduling combined with asynchronous execution provides additional performance gains beyond Asteroid’s capabilities.
@@
4.3.2  Cost Time. As illustrated in Figure 11b, Alpa, Metis, and Asteroid exhibit substantially longer parallelism strategy search time compared to DeepCEE. Alpa requires 286.3 seconds for ResNet101 and 130 seconds for GPT-2, primarily due to its exhaustive search across all possible device combinations, layer mappings, and hyperparameters (e.g., batch/micro-batch sizes), resulting in an $O(n!\cdot n!\cdot n^{2})$ search space. Metis reduces this overhead to 230 seconds (ResNet101) and 43 seconds (GPT-2) by pre-grouping devices into performance, balanced device groups and applying search-space pruning, though it still maintains $O(n!\cdot n^{3})$ time complexity. Asteroid further shortens the search time to 197 seconds (ResNet101) and 59 seconds (GPT-2) but remains slower than DeepCEE because it (1) operates on the full device set without pre-grouping and (2) incurs additional latency from runtime memory analysis and parallel strategy evaluation.
In contrast, DeepCEE achieves significantly faster strategy discovery through three key optimizations: (1) topology-aware device grouping that reduces the search space to near-constant complexity, (2) rapid candidate generation via predefined partitioning points, and (3) lightweight refinement using beam search.
4.3.3  Profit of Second-Level Computing Device Group. In Figure 11a, we evaluate the training performance when turning off DeepCEE’s model split among second-level device groups (No-Second). The experimental results show that the training performance of No-Second declines by 10.93% and 12.04% compared to DeepCEE when using ResNet101 and GPT-2, respectively. This is due to the inability to fully leverage the resources of the heterogeneous computing devices when model splitting is disabled among second-level device groups. However, its training performance is still higher than existing parallel training frameworks such as Alpa, Metis, and Hetpipe. As shown in Figure 11b, turning off DeepCEE’s second-level device group search can reduce its search time of the parallel strategy by roughly 20%.
4.4. Evaluating DeepCEE’s Adaptation